An annuity pays quarterly for 16 years at the beginning of each quarter. Payments during the first year are 250/quarter. Year 2 payments are each 105. The payments go up by 150 for each subsequent year but are level within each year. AEIR = 5%, please find the accumulated value of this annuity just prior to the payment at t=15.

Please show all steps, I don't need graphs so do not use excel, I need steps.

A.

79,798.78

B.

The answer does not appear here

C.

72,848.29

D.

154,014.68

E.

156,104.12
Answer #1

Do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

I have solved this question previously. I am attaching the steps to be followed for solving this problem.

Also, for calculation of steps in payments for around 60 times, I think it is inherent to use excel for calcualtion as well as display purposes.

Initially we try to calcuate the rate of interest which is given by

r = (1 + AEIR)(PAYMENT FREQUENCY)

r = (1+ 5%)1/4

r = 1.2272234%

Next we calculate the annuity for each payment time period using the following table where

FV = Cash flows ( 1 +r)n

Here cash flows are 250 and 105 for first 8 quarters and are stepped up by 150 in the payments after the 8th Quarter.

At t= 15 ( 15*4) = 60 Payments, we get the following table,

Time = t (60 - t) Cash flows Future Value 250 60 519.7320316 250 513.4310852 59 58 250 507.206528 3 57 250 501.0574339 56 105

.

.

.

.

48 1755 12 2031.631865 2007.001473 1982.669687 49 11 1755 10 50 1755 1958.632886 51 1755 52 2100.262493 1905 53 1905 2074.800

We find that option A is the answer.

Cheers!!!!